how much pure water must be mixed with 4 pints of 80% developer to produce a mixture that is 23% developer?

Answer by stanbon(75887)

how much pure water must be mixed with 4 pints of 80% developer to produce a mixture that is 23% developer?
------
Equation:
water + water = water
0.20*4 + 1.00x = 0.77*(4+x)
----
Multiply thru by 100 to get:
20*4 + 100x = 77*4+77*x
23x = 57*4
---
x = (57*4)/23
---
x = 9.91.. pints of water are needed
============================================
cheers,
Stan H.